# Is the Cure Still Worse Than the Disease? Test Overfitting by LLMs in Automated Program Repair

### v1

@@ Abstract.
Automated program repair has been shown to be susceptible to generating repaired code that passes on seen tests but fails on a hold-out set of hidden tests. This problem, dubbed test overfitting, has been identified and studied before the rise of large language models. We experimentally study how much test overfitting is still a problem today, using repository-level SWE-bench tasks.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Automated program repair (APR), which tries to automatically fix bugs in code, has long been an active research topic in the software engineering community (legoues_pradel_roychoudhury_2019). Thanks to SWE-bench (jimenez_et_al_2024), a recent benchmark for issue-driven repository-level APR, it is now also a hot research topic in the artificial intelligence (AI) community. Most solutions use large language models (LLMs) in prompted workflows without fine-tuning, such as SWE-Agent (yang_et_al_2024), Agentless (xia_et_al_2025), or CodeMonkeys (ehrlich_et_al_2025). SWE-bench evaluates a candidate code repair using hidden (black-box) tests. Therefore, solutions often leverage a disjoint set of seen (white-box) tests to select or improve a good candidate code repair (yang_et_al_2024; xia_et_al_2025; ehrlich_et_al_2025). Unfortunately, doing so creates the risk of test overfitting, where the code generated by APR passes white-box tests but fails black-box tests. Even though Smith et al. identified the test overfitting problem ten years ago (smith_et_al_2015), current LLM-driven APR solutions do not account for it.
Therefore, this paper sets out to answer the question: is test overfitting still a problem in the era of LLM-based APR workflows? Besides Smith et al.’s work (smith_et_al_2015) which predates LLMs, the closest related work on test overfitting is by Stroebl et al. (stroebl_kapoor_narayanan_2024). However, unlike our work, it does not use repository-level benchmarks, and uses white-box tests only to select one among a set of candidate code repairs, not to improve candidate code repairs. Some recent workflows have sophisticated loops of improving code repairs along with white-box tests (ehrlich_et_al_2025; li_et_al_2025_sstar). This approach might exacerbate test overfitting, which the papers unfortunately do not explore.
In contrast to the prior work, our paper directly investigates test overfitting in repository-level tasks with state-of-the-art (SoTA) LLM-based workflows. It starts from an initial candidate code generated by the SoTA APR solution Agentless (xia_et_al_2025) and an initial candidate test generated by the SoTA issue reproduction solution e-Otter++ (ahmed_et_al_2025_eotter). This paper compares the behavior of APR code on both white-box and black-box tests to answer the following research questions: RQ1. Can LLMs overfit on white-box tests? (Yes, they can and do.) RQ2. How does improving code based on white-box tests affect test overfitting? (Overfitting goes up for affected instances.) RQ3. How much would revealing the black-box tests increase apparent resolution rate as measured by the black-box tests? (Not vastly more than using white-box tests.) The experiments are based on the TDD-bench Verified (ahmed_et_al_2024) benchmark for test-driven development and use two LLMs, Claude-3.7 Sonnet and GPT-4o. This paper makes the following contributions.
@@
  • It measures the degree of test overfitting by LLMs in repository-level APR tasks.
@@
  • It describes techniques to reduce overfitting and reports how much they help overfitting and hurt performance.
@@
  • It also includes a limit study of how much an oracle with access to the golden black-box tests could help performance.
Our paper title is inspired by Smith et al. (smith_et_al_2015). If the disease is code bugs and the cure is APR with white-box tests, Smith et al. found that the cure was indeed worse than the disease: APR decreased performance on black-box tests. Our findings are less drastic: on average across all benchmark instances, APR with white-box tests improves performance on black-box tests. But there is also still test overfitting to beware of—the cure can still cause harm.
@@ 2. Background and Related Work
Test overfitting happens when an APR workflow generates code that passes white-box tests but fails held-out black-box tests (smith_et_al_2015). Test overfitting is similar to overfitting a model to its training data, where the model performs worse on held-out test data; but note that in this work, we use only pre-trained LLMs without fine-tuning. Instead, test overfitting means that generated repaired code is too narrow to generalize to held-out black-box tests. Stroebl et al. report that this happens when a loop repeatedly generates code for the HumanEval or MBPP benchmarks until finding code that passes white-box tests (stroebl_kapoor_narayanan_2024). Unlike Stroebl et al., we experiment with repository-level benchmarks and we also explore workflows that improve code based on tests.
SWE-bench is an APR benchmark where the input comprises an issue and the original code of an entire repository before issue resolution and the output is repaired code that must pass a held-out test set (jimenez_et_al_2024). Those held-out tests are unavailable, hidden from the APR workflow. SWE-bench Verified is a subset of 500 high-quality SWE-bench instances, targeted by most LLM-based APR workflows (chowdhury_et_al_2024). SWT-bench (mundler_et_al_2024) and TDD-Bench Verified (ahmed_et_al_2024) are reproduction test benchmarks derived from SWE-bench and SWE-Bench Verified. A generated output test must fail on the original code to reproduce the issue, and must pass on the hidden ground-truth code patch to confirm that it resolves the issue (mundler_et_al_2024).
Several LLM-based workflows leverage white-box tests. CodeT generates many code candidates and many candidate white-box tests, then uses a dual execution agreement to pick a single output code (chen_et_al_2023). Agentless generates up to 40 candidate codes for SWE-bench and generates a white-box reproduction test, then uses the test to filter the candidate codes as part of code selection (xia_et_al_2025). Both S* (li_et_al_2025_sstar) and CodeMonkeys (ehrlich_et_al_2025) also generate both candidate codes and candidate tests, followed by a loop where the LLM improves the code based on the test and vice versa. These papers might suffer from test overfitting but do not study it.
Test overfitting is related to reward hacking, where a reward is a reinforcement-learning (RL) objective, and a hack finds a loop-hole for improving the reward while circumventing its intention. CURE uses RL to co-evolve an LLM to generate both code and white-box tests, and might be susceptible to test overfitting and reward hacking (wang_et_al_2025). Baker et al. found that RL to improve frontier LLMs for repository-level APR tasks can cause the reward hack of disabling tests instead of repairing code (baker_et_al_2025).
@@ 3. Methodology
Let $c_{\mathrm{old}}$ be the original code in the repository before a given issue is fixed. This work uses Agentless (xia_et_al_2025) to generate new code $c_{\mathrm{new}}$ for APR and e-Otter++ (ahmed_et_al_2025_eotter) to generate a white-box test $t$ for issue reproduction. Being generated by LLMs, both $c_{\mathrm{new}}$ and $t$ are possibly imperfect. We run the generated test $t$ on the generated code $c_{\mathrm{new}}$. If it fails, we use a test-based code improvement approach to generate a better $c_{\mathrm{new}}$ that passes the (possibly also improved) white-box test $t$. This section briefly discusses the generation of the initial $c_{\mathrm{new}}$ and $t$ and our code improvement approach.
3.1. Code Patch Generation: Agentless
Agentless performs hierarchical localization followed by repair and patch validation (xia_et_al_2025). Localization starts by identifying suspicious files using both LLM prompting and embedding-based retrieval, filtering out irrelevant files. It then narrows down to relevant classes or functions using a concise skeleton representation of files. Finally, it pinpoints exact edit locations at the line level. Repair generates multiple candidates for new code $c_{\textrm{new}}$ in a lightweight search-replace diff format rather than rewriting entire code blocks, which minimizes hallucinations and enhances cost-efficiency. Patch validation automatically generates reproduction tests $t$ (when none are provided) and combines them with existing regression tests in the repository to select a single $c_{\textrm{new}}$ via majority voting. This modular structure allows Agentless to remain interpretable, easily debuggable, and effective without relying on complex environment interaction or planning mechanisms typical of agents. With the Claude-3.7-Sonnet model, Agentless achieves a 50.8% issue resolution rate on SWE-bench Verified.
3.2. Reproduction Test Generation: e-Otter++
e-Otter++ (ahmed_et_al_2025_eotter)is a system designed to generate reproduction tests $t$ from software issue descriptions, such as bug reports or feature requests, before the corresponding code patch exists. It supports test-driven development workflows and helps validate code patches produced by APR workflows. The system augments LLM outputs with checks and repairs, such as linter checks and fixing import statements. It also employs a self-reflective action planner that iteratively decides which parts of the code to ‘read’ and where and how to write or modify test code. After generating candidate tests, e-Otter++ obtains execution feedback, such as runtime errors or successes, and uses this information to refine prompts and guide test selection. Additionally, e-Otter++ leverages issue morphing–transforming a bug report into multiple variants to generate diverse tests. It then uses a ranker that combines running on imperfect surrogate code patches and coverage to select a single top-ranked test $t$. As a result, e-Otter++ achieves a fail-to-pass success rate of 63% on the TDD-Bench Verified benchmark using Claude-3.7-Sonnet.
3.3. Test-based Code Refinement
Figure 1. Overview of test-based code improvement.
@@ 3.3. Test-based Code Refinement
Figure 1 shows our test-based code refinement approach. We start with the initial candidate code $c_{\textrm{new}}$ from Agentless and the initial candidate test $t$ from e-Otter++. We execute the test on $c_{\textrm{old}}$ and $c_{\textrm{new}}$. If the test goes from fail-to-pass, we immediately stop and declare success. Otherwise, we collect the execution logs from both test executions, the focal functions modified by $c_{\textrm{new}}$, and the test function $t$ and provide them to an LLM-based critic. We ask the critic which part to modify (focal or test). Inspired by e-Otter++, we also collect additional information such as the buggy line, relevant issue line, and lookup function (the function the model wants to see to fix it). Then, we ask the LLM to modify the focal or test, exposing the counterpart (e.g., for focal, the white-box test) and additional information. After updating the focal or test, if the test is still not fail-to-pass, we need to decide whether to stick with our initial focal or test or update it with the recently generated focal or test, even though the test is not fail-to-pass. To do this, we use a reward function to guide our decision: $\mathit{Reward}(c_{\mathit{old}},c_{\mathit{new}},t)\,=$
$\frac{1}{3}\mathit{isFail}(t,c_{\mathrm{old}})+\frac{1}{3}\mathit{isPass}(t,c_{\mathrm{new}})+\frac{1}{3}\mathit{coverage}(c_{\mathit{old}},c_{\mathit{new}},t)$
@@
In the reward function, the first two components are binary, but coverage can vary from 0 to 1. Since we know both the patch and the test, we can compute how many deleted lines in $c_{\mathrm{old}}$ and how many added lines in $c_{\mathrm{new}}$ have been covered by the test. We divide the covered lines by the total number of added or deleted lines to compute coverage. If our new code+test pair achieves a higher score than the previous one, we replace it with the new one. We repeat the improvement loop for at most 15 iterations if the white-box test does not change from fail to pass.
@@ 4. Results
4.1. Expermental Setup
The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests. Agentless was evaluated on SWE-bench Verified, and e-Otter++ was evaluated on TDD-Bench Verified, a subset of 449 instances from SWE-bench Verified. We evaluated on all 449 instances. The experiments use two frontier models: Claude-3.7-Sonnet and GPT-4o. Our primary metric is test overfitting rate. A code patch is overfitting tests if it passes white-box tests but fails black-box tests. Dividing the number of instances where that happes by the total number of instances yields the overfitting rate.
@@ 4.1. Expermental Setup
Table 1. Test overfitting rate of automatic program repair.
@@
4.2. LLMs Overfit on White-box Tests (RQ1)
To answer this research question, we run the e-Otter++ generated white-box tests on Agentless patches and see how many of them go from fail-to-pass. Figure 2 shows that with the Claude-3.7-Sonnet model, 229 tests actually pass on the initial patches. Does this mean that the black-box golden tests will pass on these patches? The answer is no. There are 50 instances where the black-box test could not pass, signaling potential overfitting (50/229=21.8%). For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).

### v2

# Investigating Test Overfitting on SWE-bench

@@ Abstract.
Tests can be useful towards resolving issues on code repositories. However, relying too much on tests for issue resolution can lead to code that technically passes observed tests but actually misses important cases or even breaks functionality. This problem, called test overfitting, is exacerbated by the fact that issues usually lack readily executable tests. Instead, several issue resolution systems use tests auto-generated from issues, which may be imperfect. Some systems even iteratively refine code and tests jointly. This paper presents the first empirical study of test overfitting in this setting.
@@ 1. Introduction
Issue resolution systems, which automatically resolve bug reports or feature requests in code repositories, are useful. They are also challenging to build, because issue descriptions tend to be informal and code repositories are often large and complex (Jimenez et al., 2024). Several issue resolution systems leverage tests: for code candidate selection in parallel inference scaling (Arora et al., 2024; Gao et al., 2025; Jain et al., 2025; Ruan et al., 2024; Xia et al., 2025; Zhang et al., 2025), code refinement in sequential inference scaling (Chen et al., 2024; Li et al., 2025b; Yang et al., 2024; Zhang et al., 2024), or both (Ehrlich et al., 2025; Li et al., 2025c). These systems rely upon the implicit premise that tests offer a reliable signal in this setting. Unfortunately, little is known about whether this premise holds. To the contrary, startling results in adjacent settings cast doubt upon it (Baker et al., 2025; Smith et al., 2015; Stroebl et al., 2024). Relying upon tests for issue resolution can cause test overfitting: generated code that narrowly passes the observed tests but breaks other functionality. This paper empirically investigates the problem of test overfitting in repository-level issue resolution, on benchmarks such as SWE-bench (Jimenez et al., 2024).
An issue resolution system takes an issue description $d_{\textrm{issue}}$ along with the original code $c_{\textrm{old}}$ of a repository as input and produces modified code $c_{\textrm{new}}$ as output. While many repositories come with regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$, those only test existing behavior, not the desired functionality of $d_{\textrm{issue}}$. Instead, issue resolution systems are evaluated by running acceptance tests $t_{\textrm{gold}}$ on $c_{\textrm{new}}$. In SWE-bench, $t_{\textrm{gold}}$ is mined from issue-resolving pull requests (Jimenez et al., 2024). Since most open issues lack acceptance tests, $t_{\textrm{gold}}$ is not yet available during issue resolution — neither in practice nor in the benchmark. Therefore, several issue resolution systems instead generate surrogate tests $t_{\textrm{gen}}$, based only on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$. Since generated tests $t_{\textrm{gen}}$ are imperfect, over-relying on them risks test overfitting, where $c_{\textrm{new}}$ passes $t_{\textrm{gen}}$ but fails the hidden $t_{\textrm{gold}}$.
issue description $d_{\textrm{issue}}$
1JSONField values are displayed as dict when readonly in the admin.
2For example, {"foo": "bar"} would be displayed as {’foo’: bar’}.
code patch $c_{\textrm{new}}$ without refinement
1+ elif isinstance(field, JSONField):
2+ return field.prepare_value(value)
observed generated test $t_{\textrm{gen}}$
1value, disp = {"foo": "bar"}, {"foo": "bar"}’
2self.assertEqual(
3 display_for_field(value, JSONField(), empty), disp)
code patch $c^{\prime}_{\textrm{new}}$ with refinement
1+ elif isinstance(field, JSONField):
2+ try:
3+ return json.dumps(value)
4+ except (TypeError, ValueError):
5+ return empty_value_display
hidden golden test $t_{\textrm{gold}}$
1for value, disp in [
2 ({a’: {b’: c’}}, {"a": {"b": "c"}}’),
3 ([a’, b’], ["a", "b"]’),
4 (a’, "a"’),
5 ({(a’, b’): c’}, "{(’a’, ’b’): ’c’}"), # Invalid JSON.
6]:
7 self.assertEqual(
8 display_for_field(value, JSONField(), empty), disp)
Figure 1. Test overfitting example on django__django-12308.
Figure 1 shows a concrete example. Based on the issue description $d_{\textrm{issue}}$, an issue resolution system generates initial new code $c_{\textrm{new}}$ and a reproduction test generator generates a test $t_{\textrm{gen}}$. However, $c_{\textrm{new}}$ fails on $t_{\textrm{gen}}$, triggering code refinement based on observed test $t_{\textrm{gen}}$. The refined code $c^{\prime}_{\textrm{new}}$ passes $t_{\textrm{gen}}$ but overfits to it, since it fails other cases covered by the hidden golden test $t_{\textrm{gold}}$.
This paper empirically investigates test overfitting in issue resolution using popular large-scale open-source repositories and issues. It was inspired by two earlier studies. The paper “Is the cure worse than the disease?” (Smith et al., 2015) demonstrated that test overfitting is a serious problem. But unlike our paper, it predated large language models (LLMs) and focused on student projects. The “Inference scaling FLaws” (Stroebl et al., 2024) paper found test overfitting to be a limiting factor for inference scaling in LLM-based code generation. But unlike our paper, it did not study repository-level tasks, nor did it explore test-based code refinement. Overall, while these papers highlighted the problem in adjacent settings, to our knowledge, ours is the first to explore it in the setting of LLM-based issue resolution.
Our empirical study answers three research questions:
  • RQ1.
    Do LLM-generated code patches $c_{\textrm{new}}$ overfit to tests $t_{\textrm{gen}}$ generated using the same LLM?
  • RQ2.
    How does refining code $c_{\textrm{new}}$ based on observed tests $t_{\textrm{gen}}$ affect test overfitting?
  • RQ3.
    If the hidden tests $t_{\textrm{gold}}$ were revealed, would refining code based on them break other functionality?
Our experiments use two LLMs, Claude-3.7 Sonnet and GPT-4o. We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2025). We extend these systems with a test-based code refinement loop that is representative of those found in recent work (Ehrlich et al., 2025; Li et al., 2025c, b, a). The experiments are based on the TDD-bench Verified (Ahmed et al., 2024) benchmark for test-driven development, which is derived from SWE-bench (Jimenez et al., 2024). TDD-bench Verified comprises challenging real-world Python issue resolution tasks. This paper makes the following contributions.
@@
  • It measures the degree of test overfitting by LLMs in repository-level issue resolution tasks.
@@
  • It describes techniques to reduce overfitting and reports how they impact both overfitting and performance.
@@
  • It also includes a limit study of how much an oracle with access to the hidden golden tests could help performance.
This paper demonstrates that test overfitting is a clear problem in modern LLM-based issue resolution systems. And unfortunately, while attempts to mitigate this overfitting do work to some extent, they also reduce effectiveness on the original task of issue resolution. Even if golden tests were available, refining code to pass them can still break other functionality. Overall, we caution against over-reliance on tests during code generation and hope to inspire future work in reducing test overfitting.
@@ 2. Background and Related Work
Test overfitting happens when an system generates code that passes observed tests but fails held-out hidden tests (Smith et al., 2015). Test overfitting is similar to overfitting a model to its training data, where the model performs worse on held-out test data; but note that in this work, we use only pre-trained LLMs without fine-tuning. Instead, test overfitting means that generated code $c_{\textrm{new}}$ is too narrow to generalize to held-out tests $t_{\textrm{gold}}$. Stroebl et al. report that this happens when a loop repeatedly generates code for the HumanEval or MBPP benchmarks until finding code that passes observed tests (Stroebl et al., 2024). Unlike Stroebl et al., we experiment with repository-level benchmarks and we also explore test-based code refinement loops.
SWE-bench is a benchmark where the input comprises an issue description $d_{\textrm{issue}}$ and the original code $c_{\textrm{old}}$ of an entire repository before issue resolution and the output is repaired code $c_{\textrm{new}}$ (Jimenez et al., 2024). Solutions are evaluated on whether $c_{\textrm{new}}$ passes a held-out test set $t_{\textrm{gold}}$. To emulate the real-world situation where issues do not come with readily executable tests, the benchmark does not make those held-out tests available to the issue resolution system. SWE-bench Verified is a subset of 500 high-quality SWE-bench instances, targeted by most issue resolution systems (Chowdhury et al., 2024). SWT-bench (Mündler et al., 2024) and TDD-Bench Verified (Ahmed et al., 2024) are reproduction test benchmarks derived from SWE-bench and SWE-Bench Verified.
Several issue resolution systems leverage generated tests $t_{\textrm{gen}}$. CodeT generates many code candidates and many candidate tests, then uses a dual execution agreement to pick a single output code (Chen et al., 2023). Agentless generates up to 40 candidate codes for SWE-bench and generates a reproduction test, then uses the test to filter the candidate codes as part of code selection (Xia et al., 2025). S* (Li et al., 2025a), CodeMonkeys (Ehrlich et al., 2025), and InfCode (Li et al., 2025c) also generate both candidate codes and candidate tests, followed by a loop where the LLM improves the code based on the test and vice versa. These papers might suffer from test overfitting but do not study it.
Test overfitting is related to reward hacking, where a reward is a reinforcement-learning (RL) objective, and a hack finds a loop-hole for improving the reward while circumventing its intention. CURE uses RL to co-evolve an LLM to generate both code and white-box tests, and might be susceptible to test overfitting and reward hacking (Wang et al., 2025). Baker et al. found that RL to improve frontier LLMs for repository-level coding tasks can cause the reward hack of disabling tests instead of improving code (Baker et al., 2025).
@@ 3. Methodology
We assume a simple three-step pipeline for issue resolution.
  1. (1)
    Run an off-the-shelf issue resolver (Agentless (Xia et al., 2025)) on $d_{\textrm{issue}}$ and original code $c_{\textrm{old}}$ to obtain initial candidate code $c_{\textrm{new}}$.
  2. (2)
    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2025)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
  3. (3)
    Run test-based code refinement (described below) on the initial $c_{\textrm{new}}$ and $t_{\textrm{gen}}$ to obtain final candidate code $c^{\prime}_{\textrm{new}}$.
After that, we can quantify test overfitting by checking which code ($c_{\textrm{new}}$ or $c^{\prime}_{\textrm{new}}$) passes which tests ($t_{\textrm{gen}}$, $t_{\textrm{gold}}$, or $t_{\textrm{old}}$).
Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2025) have been described in prior work, here we only briefly discuss characteristics salient to our paper. First, both Agentless and e-Otter++ internally employ some parallel inference scaling. In our pipeline, we only take the single final result selected by each system and feed it into our own test-based code refinement stage, which can be viewed as a form of sequential inference scaling. Second, both Agentless and e-Otter++ internally make use of regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$ from the original code of the repository. This is a realistic reflection of how a human software engineer might consult regression tests when tackling an issue. At the same time, it only yields limited benefit, since existing tests $t_{\textrm{old}}$ tend not to test the open issue $d_{\textrm{issue}}$.
Figure 2. Overview of test-based code refinement.
@@
Figure 2 shows our test-based code refinement approach. It starts with the initial candidate code $c_{\textrm{new}}$ from Agentless and the initial candidate test $t_{\textrm{gen}}$ from e-Otter++. We execute the test on $c_{\textrm{old}}$ and $c_{\textrm{new}}$. If the test goes from fail-to-pass, we immediately stop and declare success. Otherwise, we collect the execution logs from both test executions, the focal functions modified by $c_{\textrm{new}}$, and the test function $t_{\textrm{gen}}$ and provide them to an LLM-based critic. We ask the critic which part to modify (focal or test). Inspired by e-Otter++, we also collect additional information such as the buggy line, relevant issue line, and lookup function (the function the model wants to see to fix it). Then, we ask the LLM to modify the focal code or test, exposing the counterpart (e.g., for focal, the test) and additional information. After updating the code $c_{\textrm{new}}$ or test $t_{\textrm{gen}}$, if the test is still not fail-to-pass, we need to decide whether to stick with our initial code or test or update it with the recently generated code or test, even though the test is not fail-to-pass. To do this, we use a reward function to guide our decision: $\mathit{Reward}(c_{\mathit{old}},c_{\mathit{new}},t_{\textrm{gen}})\,=$
$\frac{1}{3}\mathit{isFail}(t,c_{\mathrm{old}})+\frac{1}{3}\mathit{isPass}(t,c_{\mathrm{new}})+\frac{1}{3}\mathit{coverage}(c_{\mathit{old}},c_{\mathit{new}},t_{\textrm{gen}})$
@@
In the reward function, the first two components are binary, but coverage can vary from 0 to 1. Since we know both the code patch and the test, we can compute how many deleted lines in $c_{\mathrm{old}}$ and how many added lines in $c_{\mathrm{new}}$ have been covered by the test. We divide the covered lines by the total number of added or deleted lines to compute coverage. If our new $\langle c_{\mathit{new}},t_{\textrm{gen}}\rangle$ pair achieves a higher reward score than the previous one, we replace it with the new one. We repeat the improvement loop for at most 15 iterations if the observed test $t_{\textrm{gen}}$ does not change from fail to pass.
@@ 4. Results
The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests. Agentless was evaluated on SWE-bench Verified, and e-Otter++ was evaluated on TDD-Bench Verified, a subset of 449 instances from SWE-bench Verified. We evaluated on all 449 instances. Our primary metric is test overfitting rate. A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{new}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$. Dividing the number of instances where that happens by the total instances yields the overfitting rate.
@@
Table 1. Test overfitting rate of automated issue resolution.
@@
4.1. Overfitting without Code Refinement (RQ1)
To quantify overfitting without refinement, we run the e-Otter++ generated tests $t_{\textrm{gen}}$ on Agentless code patches $c_{\textrm{new}}$ and see how many pass. Figure 3 shows that with the Claude-3.7-Sonnet model, for 229 instances, the initial code passes the generated tests. Does this mean that the hidden golden tests $t_{\textrm{gold}}$ will pass on these patches? The answer is no. There are 50 instances where $t_{\textrm{gold}}$ failed, rendering them unresolved according to the benchmark. That means the overfitting rate among the 229 instances is $\frac{50}{229}=21.8\%$. For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).